OpenStudy (abmon98):

Thorium 32 has a half life of 1.4*10^110 years At a particular instant the activity of a sample of thorium is 120 Bq Calculate the time taken for the activity of this sample to fall to 15Bq

OpenStudy (anonymous):

n(t) = n(0) (0.5)^t' t' = time/half-life 15/120 = (0.5)^t' log(15/120) = (t') log (0.5) solve for t' and thus for time